# Can Large Reasoning Models do Analogical Reasoning under Perceptual Uncertainty?

### v1

@@ 5.1 LRMs are stronger analogical reasoners than LLMs
Table 1: Evaluating LRMs on analogical reasoning. Full task and arithmetic accuracy (%) of different LLMs and LRMs on two analogical reasoning benchmarks, I-RAVEN and I-RAVEN-X. For each model, we report if In-Context Learning (ICL) examples of the task were added to the prompt, the number of total prompts fed into the model (some techniques, such as self-consistency and disentangled prompting require querying the model multiple times), and the number of tokens generated by the model. “Range” indicates the dynamic range of the attributes’ values. “Tok.” indicates the average number of output tokens of the model. The results for GPT-4 and Llama-3 are taken from Hersche et al. (2025).
Analogical reasoning capabilities of LRMs have not been extensively evaluated to this date. In this work, we reduce this knowledge gap by testing this new generation of systems on the well-known benchmark I-RAVEN (Hu et al., 2021), as well as its more difficult extension, I-RAVEN-X (Hersche et al., 2025). Table 1 reports the results for this first proposed evaluation. We additionally include previous results on the closed-source OpenAI GPT-4 (OpenAI et al., 2024) and the open-source Llama-3 70B (Dubey et al., 2024) from Hersche et al. (2025) to allow for a one-to-one comparison between LRMs and LLMs.
@@
Firstly, we observe that LRMs can achieve results comparable to LLMs with less engineered prompts and that they generally improve the reasoning accuracy when the level of prompt engineering is on par. o3-mini, for instance, shows no drops in accuracy on I-RAVEN-X and a $6\%$ drop on I-RAVEN compared to GPT-4 while using only $\nicefrac{{1}}{{21}}$ of the prompts. When we compare the same two models on similar prompt complexities (that is, using entangled prompting in both settings, but still retaining a $\nicefrac{{1}}{{7}}$ ratio between LRMs and LLMs due to self-consistency) o3-mini emerges as a clear winner, showing a $6.5\%$ increase in accuracy and remarkably stronger performances on arithmetic reasoning. However, this comes at a cost, as shown by the number of output tokens produced by the models during inference, which is two orders of magnitude higher on average compared to LLMs.
Secondly, the results show that LRMs are much stronger reasoners than LLMs when challenged with the longer reasoning rules and attribute ranges in I-RAVEN-X. While LLMs show a massive drop in arithmetic accuracy on I-RAVEN-X, nearing $0\%$ for comparable prompt complexity, LRMs are affected by a much smaller arithmetic degradation on average, while sometimes even improving on the overall task accuracy.
Overall, we observe that o3-mini and R1 show similar performance on this analogical reasoning task, with o3-mini excelling in the standard I-RAVEN and R1 performing better on I-RAVEN-X. The distilled version of R1, on the other hand, displays weaker results compared to the original model, especially on I-RAVEN-X. To further improve the results on o3-mini, we increased the reasoning effort from medium to high and set the maximum number of reasoning to its maximum (100,000). The accuracy on I-RAVEN improves by $6$%, whereby it stays constant on the most difficult I-RAVEN-X setting despite the increased reasoning effort (2.7×2.7$\times$2.7 × reasoning tokens). Hence, o3-mini (medium) was preferred as a cost-efficient solution for the following experiments.
@@
(c) 30 ($-10$) 0.51 - - - - - - 79.2/76.9 78.8/66.5
Table 2: Evaluating LRMs and NeSy-PAR models on analogical reasoning under perceptual uncertainty. Task and arithmetic accuracy (%) of OpenAI o3-mini, DeepSeek R1, and ARLC on I-RAVEN-X (range [0,1000]) with different numbers of confounders, from 0 (no confounders, signal-to-noise ratio (SNR)$\infty$) to 10 (SNR=$-5.23$ dB), and different attributes’ distribution smoothening (bin-smoothening strategy, with different probabilities assigned to the correct value bin $p_{L}$). Three groups of experiments are reported: a) only confounders are introduced; b) only the attributes’ distribution smoothening is applied; c) both confounders and distribution smoothing are applied, together simulating perceptual uncertainty. For LRMs, we report the number of output tokens to quantify the reasoning effort adopted on average by the model to find a solution. The signal-to-noise ratio (SNR) quantifies how strong the noise is introduced by confounders in each experiment. The results for ARLC are reported as max/mean over 5 different random seeds.
@@ 5.2 LRMs are significantly challenged by reasoning under uncertainty
The results in Section 5.1 show that LRMs can solve analogical reasoning tasks more accurately than LLMs. However, would they be capable of retaining the same robustness in scenarios where uncertainty is introduced? To answer this question, we benchmark the two LRMs on the I-RAVEN-X extension proposed in Section 3. We adopt the same methodology used in the previous experiments on I-RAVEN and I-RAVEN-X, with only minor prompting modifications when strictly necessary (e.g., to provide probability distributions for attributes’ values). The empirical results of this study are reported in Table 2.
Firstly, we observe that LRMs perform significantly worse when noise factors that simulate perceptual uncertainty are integrated into the experiments. For instance, o3-mini’s accuracy dropped by $11.2\%$ and $15.2\%$ on task and arithmetic accuracy, respectively, when evaluated with 10 additional confounding attributes. R1, on the other hand, is more robust to confounders ($5.8\%$ and $12.2\%$ drops on task and arithmetic accuracy). However, it performs much worse when the attribute values’ distributions are smoothened, losing up to $19.8\%$ of task accuracy in the harshest scenario, while o3-mini shows a much smaller degradation ($5.4\%$) in this setting.
When both the confounders and distribution smoothening are evaluated together at their maximum level, we observe a sharp drop in task accuracy for both o3-mini (to $17.0\%$) and DeepSeek R1 (to $22.8\%$), bringing them close to random chance ($12.5\%$). Moreover, we observe that for o3-mini more challenging perceptual uncertainty conditions directly translate to higher numbers of reasoning tokens ($7209$ of the base setting to $18,589$ of the combined noise experiments). This trend, however, was not observed in R1, where the number of tokens was roughly constant across the different settings. An additional experiment with high reasoning effort could only slightly increase the o3-mini’s accuracy (to $31.0\%$) at the cost of $53,596$ average reasoning tokens.
Taking a step back, we observe that the overall task accuracy drop for LRMs is considerable: o3-mini loses up to $69.6\%$ accuracy and R1 up to $57.2\%$ from the standard I-RAVEN to I-RAVEN-X with perceptual uncertainty. Testing on longer reasoning relations and larger attributes’ dynamic ranges only plays a smaller part in this ($5.6\%$ for o3-mini, and even an increase in accuracy of $2.2\%$ for R1), while perceptual uncertainty accounts for most of the actual drop in accuracy.
@@ 5.3 NeSy-PAR models are robust when reasoning under uncertainty
We extend the investigation to neuro-symbolic models based on probabilistic abductive reasoning, focusing in particular on ARLC ( Camposampiero et al. (2024), improved with the entropy regularization introduced in Section 4.2). We report some of these results in Table 2, and more extensive evaluations in Appendix D. ARLC proves to be more robust when reasoning under perceptual uncertainty compared to LRMs, showing no drop in accuracy even in extremely harsh signal-to-noise conditions due to confounders thanks to the novel entropy-based regularization (up to $-20$ dB as shown in Appendix D) and maintains its high accuracy when reasoning with smoothened attributes’ distributions. Furthermore, when evaluated on the most difficult setting (group (c) in Table 2), ARLC displays much stronger results ($88.0\%$ best accuracy compared to $23.2\%$ of the best LRM). Overall, ARLC maintains a remarkably high reasoning accuracy despite the introduction of perceptual uncertainty in the trajectory I-RAVEN $\rightarrow$ I-RAVEN-X, experiencing only a modest decline (98.6% to 88.0%), significantly outperforming LRMs. ARLC can also successfully learn the set of rules underlying I-RAVEN when trained with highly uncertain attribute distributions, as shown in Appendix D.
@@ 6 Conclusion
This work addresses a substantial limitation of pre-existing symbolic analogical reasoning benchmarks used to evaluate LLMs, i.e., their lack of support for reasoning under perceptual uncertainty. Specifically, it augments an existing benchmark based on RPM, I-RAVEN-X, with confounding attributes and smooth attributes’ distributions, that together allow to simulate an imperfect perception front-end. This benchmark is then used to evaluate the latest generation of open-domain reasoning systems, Large Reasoning Models (LRMs). Compared to LLMs, LRMs achieve improved productivity to larger reasoning relations and attribute ranges. However, LRMs are still significantly challenged by the (simulated) perceptual uncertainty, which reduces the model reasoning accuracy by $69.6\%$ and 57.4% (o3-mini and R1, respectively) On the other hand, neuro-symbolic models based on probabilistic abduction achieve more robust and accurate performance but cannot directly generalize to different domains in the same way LRMs do. Overall, our results suggest that open-domain, robust analogical reasoning models are still a mirage, and future work has to be invested to achieve this objective.
@@ Appendix A Additional details on RPM and prompting
We report some examples of the prompts used in our experiments in Tables A.3, A.4, A.5, and A.6. The prompting style for embracing CoT was inspired by Wüst et al. (2024). For automatic retrieval of the model’s answer, we prompt it to provide its answer in the format “My Answer: Answer #<your answer>”. By default, answer panel #0 is predicted if no answer can be retrieved.
@@ Appendix D Additional experimental results for ARLC
On the other hand, evaluating models trained on clean data (that learned a good set of rules) with smoothed attributes’ distributions sensibly degrades the test accuracy, especially for larger values of $\sigma$. Unfortunately, not much can be done to address this. However, this is still an interesting result, as it underlines the importance of a confident front-end perception in abstract reasoning, showing that excessively flattened attributes’ distribution can seriously undermine the accuracy of the reasoning process. Finally, we observe that training with smoothened distributions generally yields better results (at least in the maximum test accuracy) than training on clean data and evaluating with smoothened distributions. This might suggest that, sometimes, integrating uncertainty in the training process can increase the robustness of the model at inference time and guarantee better performances compared to models that have trained exclusively on clean data.

### v2

@@ 5.1 LRMs are stronger analogical reasoners than LLMs
Table 1: Evaluating LRMs on analogical reasoning. Full task accuracy (i.e, % of test examples correctly predicted) and arithmetic accuracy (i.e., % of the attributes in the test examples governed by an arithmetic relation correctly predicted) of LLMs and LRMs on I-RAVEN and I-RAVEN-X. We report if In-Context Learning (ICL) examples of the task were added to the prompt, the number of total prompts fed into the model (some techniques, such as self-consistency and disentangled prompting require querying the model multiple times), and the number of tokens generated by the model. “Range” indicates the dynamic range of the attributes’ values. “Tok.” indicates the average number of output tokens of the model. The results for GPT-4 and Llama-3 are taken from Hersche et al. (2025).
@@
Analogical reasoning capabilities of LRMs have not been extensively evaluated to date. In this work, we reduce this knowledge gap by testing this new generation of systems on the well-known benchmark I-RAVEN (Hu et al., 2021), as well as its more difficult extension, I-RAVEN-X (Hersche et al., 2025). Table 1 reports the results for this first proposed evaluation. We additionally include previous results on the closed-source OpenAI GPT-4 (OpenAI et al., 2024) and the open-source Llama-3 70B (Dubey et al., 2024) from Hersche et al. (2025) to allow for a one-to-one comparison between LRMs and LLMs. Firstly, we observe that LRMs can achieve results comparable to LLMs with less engineered prompts and that they generally improve the reasoning accuracy when the level of prompt engineering is on par. o3-mini, for instance, shows no drops in accuracy on I-RAVEN-X and a $6\%$ drop on I-RAVEN compared to GPT-4 while using only $\nicefrac{{1}}{{21}}$ of the prompts. When we compare the same two models on similar prompt complexities (that is, using entangled prompting in both settings, but still retaining a $\nicefrac{{1}}{{7}}$ ratio between LRMs and LLMs due to self-consistency) o3-mini emerges as a clear winner, showing a $6.5\%$ increase in accuracy and remarkably stronger performances on arithmetic reasoning. However, this comes at a cost, as shown by the number of output tokens produced by the models during inference, which is two orders of magnitude higher on average compared to LLMs. Secondly, the results show that LRMs are much stronger reasoners than LLMs when challenged with the longer reasoning rules and attribute ranges in I-RAVEN-X. While LLMs show a massive drop in arithmetic accuracy on I-RAVEN-X, nearing $0\%$ for comparable prompt complexity, LRMs are affected by a much smaller arithmetic degradation on average, while sometimes even improving on the overall task accuracy. Overall, we observe that o3-mini and R1 show similar performance on this analogical reasoning task, with o3-mini excelling in the standard I-RAVEN and R1 performing better on I-RAVEN-X. The distilled version of R1, on the other hand, displays weaker results compared to the original model, especially on I-RAVEN-X. To further improve the results on o3-mini, we increased the reasoning effort from medium to high and set the maximum number of reasoning to its maximum (100,000). The accuracy on I-RAVEN improves by $6$%, whereby it stays constant on the most difficult I-RAVEN-X setting despite the increased reasoning effort (2.7×2.7$\times$2.7 × reasoning tokens). Hence, o3-mini (medium) was preferred as a cost-efficient solution for the following experiments.
@@
20 ($-8.24$) 0.51 - - - - - - 83.2/76.9 81.7/62.0
30 ($-10$) 0.51 - - - - - - 79.2/75.6 78.8/61.5
Table 2: Evaluating LRMs and NeSy-PAR models on analogical reasoning under simulated perceptual uncertainty. Task and arithmetic accuracy (%) on I-RAVEN-X (range [0,1000]) with different numbers of confounders, from 0 (no confounders, SNR=$\infty$) to 10 (SNR=$-5.23$ dB), and different attributes’ distribution smoothening (bin-smoothening strategy, with different probabilities assigned to the correct value bin $p_{L}$). We show experiments with: a) only confounders ; b) only the attributes’ distribution smoothing ; c) both confounders and distribution smoothing. For LRMs, we report the number of output tokens to quantify the reasoning effort adopted on average by the model to find a solution. The results for ARLC are reported as max/mean over 5 different random seeds.
@@ 5.2 LRMs are significantly challenged by reasoning under uncertainty
The results in Section 5.1 show that LRMs can solve analogical reasoning tasks more accurately than LLMs. However, would they be capable of retaining the same robustness in scenarios where uncertainty is introduced? To answer this question, we benchmark the two LRMs on the I-RAVEN-X extension proposed in Section 3. We adopt the same methodology used in the previous experiments on I-RAVEN and I-RAVEN-X, with only minor prompting modifications when strictly necessary (e.g., to provide probability distributions for attributes’ values). The empirical results of this study are reported in Table 2. Firstly, we observe that LRMs perform significantly worse when noise factors that simulate perceptual uncertainty are integrated into the experiments. For instance, o3-mini’s accuracy dropped by $11.2\%$ and $15.2\%$ on task and arithmetic accuracy, respectively, when evaluated with 10 additional confounding attributes. R1, on the other hand, is more robust to confounders ($5.8\%$ and $12.2\%$ drops on task and arithmetic accuracy). However, it performs much worse when the attribute values’ distributions are smoothened, losing up to $19.8\%$ of task accuracy in the harshest scenario, while o3-mini shows a much smaller degradation ($5.4\%$) in this setting. When both the confounders and distribution smoothening are evaluated together at their maximum level, we observe a sharp drop in task accuracy for both o3-mini (to $17.0\%$) and DeepSeek R1 (to $22.8\%$), bringing them close to random chance ($12.5\%$). Moreover, we observe that for o3-mini more challenging perceptual uncertainty conditions directly translate to higher numbers of reasoning tokens ($7209$ of the base setting to $18,589$ of the combined noise experiments). This trend, however, was not observed in R1, where the number of tokens was roughly constant across the different settings. An additional experiment with high reasoning effort could only marginally increase the o3-mini’s accuracy (to $31.0\%$) at the cost of $53,596$ average reasoning tokens.
Figure 2: LRM/ARLC reasoning time vs. accuracy.
@@ 5.3 NeSy-PAR models are robust when reasoning under uncertainty
We extend the investigation to neuro-symbolic models based on probabilistic abductive reasoning, focusing in particular on ARLC Camposampiero et al. (2024), improved with the entropy regularization introduced in Section 4.2. We report some of these results in Table 2, and more extensive evaluations in Appendix D. ARLC proves to be more robust when reasoning under perceptual uncertainty compared to LRMs, showing no drop in accuracy even in extremely harsh SNR conditions due to confounders thanks to the novel entropy-based regularization (up to $-20$ dB as shown in Appendix D) and maintains its high accuracy when reasoning with smoothened attributes’ distributions. Furthermore, when evaluated on the most difficult setting (group (c) in Table 2), ARLC displays much stronger results ($88.0\%$ best accuracy compared to $23.2\%$ of the best LRM). Overall, ARLC maintains a remarkably high reasoning accuracy in the trajectory I-RAVEN $\rightarrow$ I-RAVEN-X, experiencing only a modest decline (98.6% to 88.0%) and significantly outperforming LRMs. ARLC can also successfully learn the set of rules underlying I-RAVEN when trained with highly uncertain attribute distributions, as shown in the ablation included in Appendix D. Finally, we directly compare ARLC and o3-mini on the accuracy-cost trade-off scaling behavior, as depicted in Figure 2. We use inference time as a proxy for computational cost. The results show that ARLC consistently defines the Pareto frontier, highlighting a more favorable scaling profile compared to o3-mini. See Appendix E for more details.
6 Discussion and limitations
Overall, we can draw different conclusions from the experimental results reported in Section 5. First, we confirm that LRMs yield substantial improvements over LLMs also in analogical reasoning tasks, as demonstrated in Section 5.1. Notably, arithmetic reasoning performance shows marked gains, with improvements reaching up to 65.4% in certain settings. However, this favorable trend shifts upon the introduction of simulated perceptual uncertainty. The overall task accuracy drop for LRMs is considerable: o3-mini loses up to $69.6\%$ accuracy and R1 up to $57.2\%$ from the standard I-RAVEN to I-RAVEN-X with uncertainty. Testing on longer reasoning relations and larger attributes’ dynamic ranges only plays a smaller part in this ($5.6\%$ for o3-mini, and even an increase in accuracy of $2.2\%$ for R1), while perceptual uncertainty accounts for most of the actual drop in accuracy. On the other hand, neuro-symbolic probabilistic abductive reasoning methods show stronger robustness and better scaling laws compared to LRMs. While prior works have circumvented the need for artificially induced perceptual uncertainty by integrating neural perception modules into the reasoning pipeline (Hersche et al., 2023; Shah et al., 2022; Ahrabian et al., 2024a; Jiang et al., 2024a), the unimodal nature of LLMs and LRMs constrains our approach to purely symbolic datasets. On the other hand, we note that the simulated perceptual uncertainty introduced in our setup may be simultaneously overly simplistic and unnecessarily complex. Specifically, the three-bin discretization strategy constitutes an oversimplification of the behavior of a real perception module, which would likely distribute uncertainty across a broader set of elements and introduce additional errors not currently modeled. At the same time, encoding this uncertainty leads to significantly longer and less efficient LRM prompts, which may in turn impair the model’s reasoning abilities. We leave a more thorough investigation of these limitations to future work
@@ 7 Conclusion
This work addresses the lack of support for reasoning under perceptual uncertainty in symbolic analogical reasoning benchmarks. Specifically, it augments an existing benchmark based on RPM, I-RAVEN-X, with confounding attributes and smooth attributes’ distributions, that together allow for the simulation of an imperfect perception front-end. This benchmark is then leveraged to evaluate the latest generation of open-domain reasoning systems, Large Reasoning Models (LRMs). Compared to LLMs, LRMs achieve improved productivity for larger reasoning relations and attribute ranges. However, LRMs are still significantly challenged by the (simulated) perceptual uncertainty. On the other hand, neuro-symbolic models based on probabilistic abduction achieve more robust and accurate performance, but cannot directly generalize to different domains in the same way LRMs do. Overall, our results suggest that open-domain, robust analogical reasoning models are still a mirage, and future work has to be invested to achieve this objective.
[…]
This work is supported by the Swiss National Science Foundation (SNF), grant 10002666.
@@ Appendix A Additional details on RPM and prompting
[…]
@@ Appendix D Additional experimental results for ARLC
On the other hand, evaluating models trained on clean data (that learned a good set of rules) with smoothed attributes’ distributions sensibly degrades the test accuracy, especially for larger values of $\sigma$. Unfortunately, not much can be done to address this. However, this is still an interesting result, as it underlines the importance of a confident front-end perception in analogical reasoning, showing that excessively flattened attributes’ distribution can seriously undermine the accuracy of the reasoning process. Finally, we observe that training with smoothened distributions generally yields better results (at least in the maximum test accuracy) than training on clean data and evaluating with smoothened distributions. This might suggest that, sometimes, integrating uncertainty in the training process can increase the robustness of the model at inference time and guarantee better performances compared to models that have trained exclusively on clean data.
Appendix E Accuracy-cost scaling behavior experiments
In this section, we include additional details on Figure 2, reporting the accuracy-cost tradeoff for different models: o3-mini with different reasoning budgets (medium and high), GTP-4o, and ARLC. We opt to quantify cost in terms of average inference time to solve a test example; some of the models (e.g., OpenAI o3-mini) are, in fact, only accessible through API, and an exact quantification of the computation resources required by the inference is not publicly available. For ARLC, the time measurements are performed accelerating the model on a NVIDIA A100-40G GPU using batch size 1.